OpenStudy (anonymous):

(r/v)-5 Express using positive exponents. Then, if possible, simplify.

OpenStudy (anonymous):

\[\left(\begin{matrix}r \\ v\end{matrix}\right)-5\]

OpenStudy (anonymous):

Did you mean\[(\frac rv) ^{-5}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(\frac{v}{r})^5\]

jimthompson5910 (jim_thompson5910):

\[\Large \left(\frac{r}{v}\right)^{-5}\] \[\Large \left(\frac{v}{r}\right)^{5}\] \[\Large \frac{v^5}{r^5}\] So \[\Large \left(\frac{r}{v}\right)^{-5}=\frac{v^5}{r^5}\]

OpenStudy (anonymous):

I don't keep the -5

jimthompson5910 (jim_thompson5910):

no, you convert that negative exponent into a positive one

OpenStudy (anonymous):

yes jim is right

OpenStudy (anonymous):

thats what i originally thought but in my text book the example showed it staying but i think i believe you better than this thing in my hand

jimthompson5910 (jim_thompson5910):

well if the negative exponent is staying, then the book ought to be leaving because that's incorrect

jimthompson5910 (jim_thompson5910):

besides, it even says "Express using positive exponents"

OpenStudy (anonymous):

jim thompson is right. I had hard time typing fractions

OpenStudy (anonymous):

okay tks a lot i wish i was good at this but has it is i truly stink lol tks again

OpenStudy (anonymous):

you are welcome. You are good at something else :-)))

jimthompson5910 (jim_thompson5910):

well 90% of it is in your head, so if you look at it in another way, who knows

OpenStudy (anonymous):

yes science, reading everything but math lol

OpenStudy (anonymous):

and i need the math for Science

OpenStudy (anonymous):

@ Jim i can confuse myself very easily when it comes to math i have to admit i do let it intimidate me a lot

jimthompson5910 (jim_thompson5910):

so all you have to do is scare it more and it won't be so bad

OpenStudy (anonymous):

lol

jimthompson5910 (jim_thompson5910):

just grrr at your paper....ok don't know how far that'll get you, but you might have fun doing it

jimthompson5910 (jim_thompson5910):

it'll definitely blow off stress and calm your nerves

jimthompson5910 (jim_thompson5910):

guess what I'm saying is that if you have fun doing it, it won't be considered work and it'll come more naturally and easier

OpenStudy (anonymous):

thanks i needed that i can't believe that its 2am here and i am still up studying a few more problems and then i am calling it a night thanks again i might be back if get stuck

jimthompson5910 (jim_thompson5910):

k I'll have the spatula ready...

OpenStudy (anonymous):

okay armed and ready lock and loaded lol

jimthompson5910 (jim_thompson5910):

let's just hope our heads don't get esploded

OpenStudy (anonymous):

lol we'll be known as empty heads

OpenStudy (anonymous):

Jim you still here?

jimthompson5910 (jim_thompson5910):

yeah what's up